# Revealing and Reducing Gender Biases in Vision and Language Assistants (VLAs)

### v1

@@ Abstract
Pre-trained large language models (LLMs) have been reliably integrated with visual input for multimodal tasks. The widespread adoption of instruction-tuned image-to-text vision-language assistants (VLAs) like LLaVA and InternVL necessitates evaluating gender biases. We study gender bias in 22 popular open-source VLAs with respect to personality traits, skills, and occupations. Our results show that VLAs replicate human biases likely present in the data, such as real-world occupational imbalances. Similarly, they tend to attribute more skills and positive personality traits to women than to men, and we see a consistent tendency to associate negative personality traits with men. To eliminate the gender bias in these models, we find that finetuning-based debiasing methods achieve the best tradeoff between debiasing and retaining performance on downstream task. We argue for pre-deploying gender bias assessment in VLAs and motivate further development of debiasing strategies to ensure equitable societal outcomes.
@@ 1 Introduction
Rapid progress in large language models (LLMs) has sparked a wave of innovation fusing visual encoding modules with LLMs, which eventually leads to vision-language models (VLMs) capable of processing both textual and visual inputs (Li et al., 2023b; Alayrac et al., 2022). With vision-language instruction finetuning, VLMs (Liu et al., 2024a; c; Zhu et al., 2024) have become assistants capable of comprehending and executing diverse task instructions. These instruction-tuned VLMs, i.e. vision-language assistants (VLAs), now have a huge potential for interacting with diverse user populations in our society. However, it is clear that social bias, especially gender bias, present in generative models can strengthen stereotypes, reinforce existing discrimination, and exacerbate gender inequalities (Bender et al., 2021; Hirota et al., 2022), which is not desirable. Therefore, identifying and mitigating biases in VLAs is essential for improving fairness, inclusivity, and their ethical deployment in our digital age.
@@
We study gender bias with respect to personality traits (Kurita et al., 2019) and workplace-related bias, which can have major real-world implications (Heilman et al., 2024). Accordingly, we design specific prompts targeting personality traits, work-related skills, and occupations. We then curate gender-balanced subsets from annotated image datasets, namely FairFace (Karkkainen & Joo, 2021), MIAP (Schumann et al., 2021), Phase (Garcia et al., 2023), and PATA (Seth et al., 2023). Images, along with prompts, are presented to the VLAs in a visual question answering (VQA) format, and responses are analyzed to discern systematic distributional differences for females and males. This analysis provides valuable insights into latent associations learned by VLAs, contributing to a deeper understanding of their shortcomings. After having discovered that VLAs are gender biased, we turn towards eliminating these biases by applying four possible debiasing methods, namely Finetuning, Prompt Tuning, Prompt Engineering, and Pruning, to a representative subset of models. Our aim is to reduce gender biases while preserving accuracy on downstream tasks.
@@
To summarize, we make the following contributions. (1) We evaluate a diverse range of 22 open-source VLAs, including large-scale models up to 34B parameters and small-scale models with 1B parameters. We thoroughly assess these 22 VLAs in terms of their susceptibility to gender bias in personality traits, work-related skills, and occupations. (2) Our analysis indicates that positive personality traits are more frequently assigned to females, while negative ones are more associated with males. Furthermore, VLAs attribute the majority of the skills evaluated in this study to women rather than to men. Regarding occupations, we find that a subset of models replicates gender imbalances found in the real world, supporting prior observations on text-to-image models and LLMs (Bianchi et al., 2023; Seshadri et al., 2024; Gorti et al., 2024). (3) Our experiments on debiasing show that finetuning yields the best tradeoff between reducing gender bias and maintaining performance on general-purpose benchmarks. Concretely, full finetuning of all parameters results in more aggressive gender debiasing while reducing performance more, and using LoRAs (Hu et al., 2022) results in less aggressive debiasing while reducing performance less.
@@ 2.2 Mitigating Dataset Bias
Dataset bias may introduce confounders, i.e. attributes of individuals or in the background that correlate with gender, where the model may have a bias regarding the attribute but not regarding gender per se. For example, a model may attribute skills more frequently to individuals wearing suits, and in this case a higher proportion of men wearing suits than women in the data may be mistakenly seen as gender bias, although the bias actually is about clothing.
@@ 3.2 Debiasing Vision and Language Assistants (DB-VLAs)
To mitigate model biases, we employ five different techniques: Full finetuning; LoRA finetuning (Hu et al., 2022); Prompt Tuning (Lester et al., 2021); Pruning and Prompt Engineering. All techniques except from Prompt Engineering use the same loss function defined below.
@@
Full Finetuning Here, we optimize all parameters in the transformer blocks of the VLA’s LLM component by gradient descent.
LoRA Finetuning Instead of optimizing all parameters in the LLMs’ transformer blocks, we train low-rank adapters (LoRAs) on all linear layers in the transformer decoder blocks. In all cases considered here, the LoRA rank is 128. Using LoRAs instead of full finetuning is memory efficient and allows dynamically switching adapters on or off.
@@
Prompt Engineering Instead of gradient-based model finetuning, we also evaluate adding manually crafted debiasing instructions to all prompts. Concretely, Howard et al. (2024) propose to add instructions such as “Please, be mindful that people should not be judged based on their race, gender, age, body type, or other physical characteristics.” before or after the original prompts. To evaluate this approach in debiasing VLAs in our scenario, we select the three suitable debiasing instructions from (Howard et al., 2024) and add them before or after the original prompts. Two main advantages of prompt engineering over optimization-based methods are that model parameters stay unchanged, and we do not need to calculate gradients, which is costly.
@@ 4.2 Evaluating Gender Bias in VLAs: An Extensive Analysis
Personality Traits Results presented in Fig. 2 (top) exhibit a clear pattern: Negative adjectives, i.e. “moody”, “arrogant”, “stubborn”, and “irritable”, are more associated with males than with females. Positive traits, however, are not attributed to one single gender. Furthermore, only one adjective that is attributed more to males than to females is positive, namely “humble”. These trends extend beyond the five most male- and female-biased traits (see Appendix E).
@@
When comparing models using our ranking criterion for biasedness defined in Section 3.1, we find that, in particular, the models in the InternVL2 series exhibit high gender-occupation bias ($>80$% of occupations with significant bias for InternVL2-4B, InternVL2-8B, and InternVL2-26B). These are followed by models from the LLaVA-1.6 series (LLaVA-1.6-Hermes-34B: 80% and LLaVA-1.6-Mistral-7B: 72%) as well as the Bunny series (Bunny-8B:L 78% and Bunny-4B: 70%). However, the remaining models in the LLaVA-1.6 series (LLaVA-1.6-Vicuna-13B and LLaVA-1.6-Vicuna-7B) are among the less biased models overall, which is likely an effect of the LLM, Vicuna in this case. The model series with the least pronounced gender-occupation bias are LLaVA-1.5 and MobileVLM ($<40\%$ of occupations with significant bias except for LLaVA-RLHF-13B: 53% and MobileVLM-7B: 55%). From this observation, we conclude that the most recent and most powerful models also have the strongest gender-occupation bias, which is surprising.
@@ 4.3 Mitigating gender biases in VLAs
The results in Table 2 indicate a tradeoff between more aggressive debiasing and performance retention on general-purpose benchmarks. Prompt Tuning reduces the measurable gender bias the most but also severely impacts the performance of models. For example, the average score on MME is merely 762 (down from 1774). On the other hand, prompt engineering does not affect downstream performance much, but it also has little effect on gender bias. For pruning, we find that higher pruning ratios result in lower gender bias scores but, as expected, hurt performance. Finally, direct finetuning seems to offer the best tradeoff. Gender bias is reduced considerably, while performance is not impacted as greatly as with other methods. Full finetuning is more detrimental to performance but reduces gender bias more while finetuning with LoRAs reduces gender bias less but also preserves performance better.
Figure 3: Comparison of Full Finetuning (middle bar) and LoRA Finetuning (right bar) as debiasing methods (original VLA = left bar in each VLA). For each prompt group and evaluated model, we show the ratio of traits/skills/occupations with significant bias for the original and debiased models.
Comparing Full finetuning and LoRA finetuning in detail on five VLAs. In Fig. 3, we provide a close-up comparison of Full Finetuning and LoRA Finetuning as the best-performing debiasing techniques. We show the ratio of personality traits, skills, and occupations before and after debiasing. Full Finetuning overall performs better than LoRA Finetuning wrt. reducing gender bias. Interestingly, in the case of LLaVA-1.5-13B, fine-tuning even increases bias, and Full Finetuning increases bias more than LoRA fine-tuning. We also note that among the five models considered here, LLaVA-1.5-13B is the least biased model to begin with. This could indicate that fine-tuning is too aggressive in already only weakly biased models. However, for other models, finetuning reduces bias consistently and, in many cases, strongly.
In particular, on InternVL2-8B, debiasing is very effective, reducing gender bias by half on the personality traits prompt group and almost completely removing it on the skills and occupations prompt groups. This is particularly relevant because InternVL2-8B is the strongest of the five compared models (Chen et al., 2024a), but also the most biased model. In conclusion, we recommend using finetuning to debias VLAs. In case more cautious finetuning is desired, LoRAs should be used, and if more aggressive debiasing is necessary, full finetuning is the better option.
@@
Qualitative results. In Fig. 4, we show examples of how debiasing yields more uniform predictions over the answer options in our prompts. For two images, we see that after debiasing by Prompt Tuning, Full Finetuning, or LoRA Finetuning, no option is predicted with high confidence. The VLA (in this case LLaVA-1.6-Vicuna-7B) claims for the image of a woman in the top row that she is creative and is well suited to be a sewer. Both answers can clearly not be deduced from the image, as there is no relevant information shown for these attributes.
@@
Similarly, the VLA confidently assumes that the man shown in the bottom row of Fig. 4 can use logical approaches. While this may be a desirable answer pragmatically, it is nonetheless not justified by any content in the image. After debiasing, “yes” still is the answer with highest probability, but the distribution over all three options is close to uniform when debiasing by Prompt Tuning, Full Finetuning, or LoRA Finetuning.
@@ 5 Related Work
We use instruction-tuned vision-language assistants (VLAs), which have demonstrated unprecedented visual understanding and reasoning abilities. The open-source VLAs included in this study are based on a LLM component, such as Vicuna (Chiang et al., 2023) or InternLM (Cai et al., 2024), and a vision component, usually a CLIP variant (Radford et al., 2021; Cherti et al., 2023; Zhai et al., 2023). These components are jointly tuned on suitable data to yield a full multimodal model (Liu et al., 2024c; Tong et al., 2024).
Gender bias in LLMs is a vast field, therefore we refer the reader to a recent study by Gallegos et al. (2024) for an overview and concentrate our literature review on vision-language models. For CLIP, Janghorbani & De Melo (2023) investigated associations between different demographics and positive/negative words, Zhou et al. (2022) found that models often associate stereotypical captions with images, Agarwal et al. (2021) found that CLIP in some cases labels humans with non-human categories, and associates especially men with criminal categories in some cases, Hall et al. (2024) demonstrated gender bias with regard to occupations, and Berg et al. (2022) demonstrated skewed associations between gendered images and positive/negative adjectives. Furthermore, (Tanjim et al., 2024; Luccioni et al., 2024; Wu et al., 2023) have studied how gender bias in CLIP transfers to text-to-image models or image editing models. Our work differs from these studies in that we analyze gender bias in VLAs, which needs a different evaluation method beyond image and text embeddings.
Among works that studied gender bias in VLAs, Sathe et al. (2024) evaluate gender-profession bias in different types of VLAs, such as T2I and I2T models, while Fraser & Kiritchenko (2024) study how well VLAs disambiguate occupation descriptions for gendered images. Ruggeri et al. (2023) studies bias based on distributional differences in the probability of answering “yes” to prompts asking for given attributes. Zhang et al. (2024); Xiao et al. (2024); Howard et al. (2024) evaluate gender bias using synthetic data, focusing on occupation-related image content (Xiao et al., 2024; Howard et al., 2024) or holistic scores to rank models for biasedness (Zhang et al., 2024). However, automatic image editing methods could introduce new artifacts, which may lead to biases. Li et al. (2024b) benchmark 10 open-source VLAs across four primary aspects (faithfulness, privacy, safety, and fairness) using free-form generation, which requires GPT-4V or human assessment for evaluation. In contrast to these studies, we study associations of gendered diverse image sets and precisely defined sets of gender bias concepts such as personality traits, skills, and occupations on 22 recent and capable VLAs unprecedented in any previous study on gender bias in VLAs. Our VQA setup also circumvents the need to evaluate free-form generations by human annotators.

### v2

@@ Abstract
Pre-trained large language models (LLMs) have been reliably integrated with visual input for multimodal tasks. The widespread adoption of instruction-tuned image-to-text vision-language assistants (VLAs) like LLaVA and InternVL necessitates evaluating gender biases. We study gender bias in 22 popular open-source VLAs with respect to personality traits, skills, and occupations. Our results show that VLAs replicate human biases likely present in the data, such as real-world occupational imbalances. Similarly, they tend to attribute more skills and positive personality traits to women than to men, and we see a consistent tendency to associate negative personality traits with men. To eliminate the gender bias in these models, we find that fine-tuning-based debiasing methods achieve the best trade-off between debiasing and retaining performance on downstream tasks. We argue for pre-deploying gender bias assessment in VLAs and motivate further development of debiasing strategies to ensure equitable societal outcomes. Code is available at https://github.com/ExplainableML/vla-gender-bias.
@@ 1 Introduction
Rapid progress in large language models (LLMs) has sparked a wave of innovation fusing visual encoding modules with LLMs, leading to vision-language models (VLMs) capable of processing both textual and visual inputs (Li et al., 2023b; Alayrac et al., 2022). With vision-language instruction fine-tuning, VLMs (Liu et al., 2024a; c; Zhu et al., 2024) have become assistants capable of comprehending and executing diverse task instructions. These instruction-tuned VLMs, i.e. vision-language assistants (VLAs), have a huge potential for interacting with diverse user populations in our society. However, social biases, especially gender bias, present in generative models can strengthen stereotypes, reinforce existing discrimination, and exacerbate gender inequalities (Bender et al., 2021; Hirota et al., 2022), which is not desirable. Therefore, identifying and mitigating biases in VLAs is essential for improving fairness, inclusivity, and their ethical deployment in our digital age.
@@
We study gender bias with respect to personality traits (Kurita et al., 2019) and workplace-related bias, which can have major real-world implications (Heilman et al., 2024). Accordingly, we design specific prompts targeting personality traits, work-related skills, and occupations. We then curate gender-balanced subsets from annotated image datasets, namely FairFace (Karkkainen & Joo, 2021), MIAP (Schumann et al., 2021), Phase (Garcia et al., 2023), and PATA (Seth et al., 2023). Images, along with prompts, are presented to the VLAs in a visual question answering (VQA) format, and responses are analyzed to discern systematic distributional differences for females and males. This analysis provides valuable insights into latent associations learned by VLAs, contributing to a deeper understanding of their shortcomings. After having discovered that VLAs are gender biased, we turn towards eliminating these biases by applying four possible debiasing methods, namely Fine-tuning, Prompt Tuning, Prompt Engineering, and Pruning, to a representative subset of models. Our aim is to reduce gender biases while preserving accuracy on downstream tasks.
@@
To summarize, we make the following contributions. (1) We evaluate a diverse range of 22 open-source VLAs, including large-scale models up to 34B parameters and small-scale models with 1B parameters. We thoroughly assess these 22 VLAs in terms of their susceptibility to gender bias in personality traits, work-related skills, and occupations. (2) Our analysis indicates that positive personality traits are more frequently assigned to females, while negative ones are more associated with males. Furthermore, VLAs attribute the majority of the skills evaluated in this study to women rather than to men. Regarding occupations, we find that a subset of models replicates gender imbalances found in the real world, supporting prior observations on text-to-image models and LLMs (Bianchi et al., 2023; Seshadri et al., 2024; Gorti et al., 2024). (3) Our experiments on debiasing show that fine-tuning yields the best trade-off between reducing gender bias and maintaining performance on general-purpose benchmarks. Concretely, full fine-tuning of all parameters results in more aggressive gender debiasing while reducing performance more, and using LoRAs (Hu et al., 2022) results in less aggressive debiasing while reducing performance less.
@@ 2.2 Mitigating Dataset Bias
Dataset bias may introduce confounders, i.e. attributes of individuals or in the background that correlate with gender, where the model may have a bias regarding the attribute but not regarding gender. For example, a model may attribute skills more frequently to individuals wearing suits, and in this case a higher proportion of men wearing suits than women in the data may be mistakenly seen as gender bias, although the bias actually is about clothing.
@@ 3.2 Debiasing Vision and Language Assistants (DB-VLAs)
To mitigate model biases, we employ five different techniques: Full fine-tuning; LoRA fine-tuning (Hu et al., 2022); Prompt Tuning (Lester et al., 2021); Pruning and Prompt Engineering. All techniques except from Prompt Engineering use the same loss function defined below.
@@
Full Fine-tuning Here, we optimize all parameters in the transformer blocks of the VLA’s LLM component by gradient descent.
LoRA Fine-tuning Instead of optimizing all parameters in the LLMs’ transformer blocks, we train low-rank adapters (LoRAs) on all linear layers in the transformer decoder blocks. In all cases considered here, the LoRA rank is 128. Using LoRAs instead of full fine-tuning is memory efficient and allows dynamically switching adapters on or off.
@@
Prompt Engineering Instead of gradient-based model fine-tuning, we also evaluate adding manually crafted debiasing instructions to all prompts. Concretely, Howard et al. (2024) propose to add instructions such as “Please, be mindful that people should not be judged based on their race, gender, age, body type, or other physical characteristics.” before or after the original prompts. To evaluate this approach in debiasing VLAs in our scenario, we select the three suitable debiasing instructions from (Howard et al., 2024) and add them before or after the original prompts. Two main advantages of prompt engineering over optimization-based methods are that model parameters stay unchanged, and we do not need to calculate gradients, which is costly.
@@ 4.2 Evaluating Gender Bias in VLAs: An Extensive Analysis
Personality Traits Results presented in Fig. 2 (top) exhibit a clear pattern: Negative adjectives, i.e. “moody”, “arrogant”, “stubborn”, and “irritable”, are more associated with males than with females. Positive traits, however, are not attributed to one single gender. For example, the positive adjective “humble” is attributed more to males than to females. More examples, in particular “wise” and “loyal” are shown in Fig. 12 in Appendix E. Thus, the trends found in Fig. 2 extend beyond the five most male- and female-biased traits.
@@
When comparing models using our ranking criterion for bias defined in Section 3.1, we find that, in particular, the models in the InternVL2 series exhibit high gender-occupation bias ($>80$% of occupations with significant bias for InternVL2-4B, InternVL2-8B, and InternVL2-26B). These are followed by models from the LLaVA-1.6 series (LLaVA-1.6-Hermes-34B: 80% and LLaVA-1.6-Mistral-7B: 72%) as well as the Bunny series (Bunny-8B:L 78% and Bunny-4B: 70%). However, the remaining models in the LLaVA-1.6 series (LLaVA-1.6-Vicuna-13B and LLaVA-1.6-Vicuna-7B) are among the less biased models overall, which is likely an effect of the LLM, Vicuna in this case. The model series with the least pronounced gender-occupation bias are LLaVA-1.5 and MobileVLM ($<40\%$ of occupations with significant bias except for LLaVA-RLHF-13B: 53% and MobileVLM-7B: 55%). From this observation, we conclude that the most recent and most powerful models also have the strongest gender-occupation bias, which is surprising.
@@ 4.3 Mitigating gender biases in VLAs
The results in Table 2 indicate a trade-off between more aggressive debiasing and performance retention on general-purpose benchmarks. Prompt Tuning reduces the measurable gender bias the most but also severely impacts the performance of models. For example, the average score on MME is merely 762 (down from 1774). On the other hand, prompt engineering does not affect downstream performance much, but it also has little effect on gender bias. For pruning, we find that higher pruning ratios result in lower gender bias scores but, as expected, hurt performance. Finally, direct fine-tuning seems to offer the best trade-off. Gender bias is reduced considerably, while performance is not impacted as greatly as with other methods. Full fine-tuning is more detrimental to performance but reduces gender bias more while fine-tuning with LoRAs reduces gender bias less but also preserves performance better.
Figure 3: Comparison of Full Fine-tuning (middle bar) and LoRA Fine-tuning (right bar) as debiasing methods (original VLA = left bar in each VLA). For each prompt group and evaluated model, we show the ratio of traits/skills/occupations with significant bias for the original and debiased models.
Comparing Full fine-tuning and LoRA fine-tuning in detail on five VLAs. In Fig. 3, we provide a close-up comparison of Full Fine-tuning and LoRA Fine-tuning as the best-performing debiasing techniques. We show the ratio of personality traits, skills, and occupations before and after debiasing. Full Fine-tuning overall performs better than LoRA Fine-tuning wrt. reducing gender bias. Interestingly, in the case of LLaVA-1.5-13B, fine-tuning even increases bias, and Full Fine-tuning increases bias more than LoRA fine-tuning. We also note that among the five models considered here, LLaVA-1.5-13B is the least biased model to begin with. This could indicate that fine-tuning is too aggressive in already only weakly biased models. However, for other models, fine-tuning reduces bias consistently and, in many cases, strongly.
In particular, on InternVL2-8B, debiasing is very effective, reducing gender bias by half on the personality traits prompt group and almost completely removing it on the skills and occupations prompt groups. This is particularly relevant because InternVL2-8B is the strongest of the five compared models (Chen et al., 2024a), but also the most biased model. In conclusion, we recommend using fine-tuning to debias VLAs. In case more cautious fine-tuning is desired, LoRAs should be used, and if more aggressive debiasing is necessary, full fine-tuning is the better option.
@@
Qualitative results. In Fig. 4, we show examples of how debiasing yields more uniform predictions over the answer options in our prompts. For two images, we see that after debiasing by Prompt Tuning, Full Fine-tuning, or LoRA Fine-tuning, no option is predicted with high confidence. The VLA (in this case LLaVA-1.6-Vicuna-7B) claims for the image of a woman in the top row that she is creative and is well suited to be a sewer. Both answers can clearly not be deduced from the image, as there is no relevant information shown for these attributes.
@@
Similarly, the VLA confidently assumes that the man shown in the bottom row of Fig. 4 can use logical approaches. While this may be a desirable answer pragmatically, it is nonetheless not justified by any content in the image. After debiasing, “yes” still is the answer with highest probability, but the distribution over all three options is close to uniform when debiasing by Prompt Tuning, Full Fine-tuning, or LoRA Fine-tuning.
@@ 5 Related Work
Gender bias in LLMs is a vast field, therefore we refer the reader to a recent study by Gallegos et al. (2024) for an overview and concentrate our literature review on vision-language models. For CLIP, Janghorbani & De Melo (2023); Zhou et al. (2022); Agarwal et al. (2021); Hall et al. (2024); Berg et al. (2022) analyzed and found gender bias. Similarly, (Tanjim et al., 2024; Luccioni et al., 2024; Wu et al., 2023) have studied how gender bias in CLIP transfers to text-to-image models or image editing models. Our work differs from these studies in that we analyze gender bias in VLAs, which needs a different evaluation method beyond image and text embeddings.
Earlier works on gender bias in vision-language models relying on masked language modeling have studied the association of gender and objects (Srinivasan & Bisk, 2022), gender and activities or occupations (Zhang et al., 2022), and performance disparity wrt. gender on different benchmarks (Cabello et al., 2023). A comprehensive review of early works on vision-and-language bias and debiasing is provided in (Lee et al., 2023).
Among works that studied gender bias in VLAs, Sathe et al. (2024) evaluate gender-profession bias in different types of VLAs, such as T2I and I2T models, while Fraser & Kiritchenko (2024) study how well VLAs disambiguate occupation descriptions for gendered images. (Wu et al., 2024) measure discrepancies in recall across different genders when predicting occupations shown in images. Ruggeri & Nozza (2023) studies bias based on distributional differences in the probability of answering “yes” to prompts asking for given attributes. Zhang et al. (2024); Xiao et al. (2024); Howard et al. (2024) evaluate gender bias using synthetic data, focusing on occupation-related image content (Xiao et al., 2024; Howard et al., 2024) or holistic scores to rank models for bias (Zhang et al., 2024). However, automatic image editing methods could introduce new artifacts, which may lead to biases. Li et al. (2024b) benchmark 10 open-source VLAs across four primary aspects (faithfulness, privacy, safety, and fairness) using free-form generation, which requires GPT-4V or human assessment for evaluation. Recently, Girrbach et al. (2024) analyzed a smaller set of VLAs for gender bias regarding skills only, using similar methods. In contrast to these studies, we study associations of gendered diverse image sets and precisely defined sets of gender bias concepts such as personality traits, skills, and occupations on 22 recent and capable VLAs unprecedented in any previous study on gender bias in VLAs. Our VQA setup also circumvents the need to evaluate free-form generations by human annotators.